I'm having trouble with quadratice equations here is a problem that I cannot seem to solve: 9x^2+9x=2=0
Answer by stanbon(75887)

I'm having trouble with quadratic equations; here is a problem that I cannot seem to solve: 9x^2+9x+2=0
Note: I put a "+" in place of the "=".
---------------
Use the quadratic formula to get:
x = [-9 +- sqrt(81-4*9*2)]/[2*9]
x = [-9 +- sqrt(9)]/18
x = [-9 +- 3] /18
x = -12/18 or x = -6/18
x = -2/3 or x = -1/3
-----------------------
Using factoring:
9x^2+9x+2=0
9x^2 + 3x + 6x + 2 = 0
3x(3x+1) + 2(3x+1) = 0
(3x+1)(3x+2) = 0
x = -1/3 or x = -2/3
==========================
Cheers,
Stan H.